Castiglianos Theorem Calculate elongation of the rod under uniaxial loading
Answer #1

media%2F144%2F14445af0-2966-4aa9-a40c-69F=P dx F-P  

Here:F-?at any r → At every section with only axial load dF -= 1 do PLo → Elongation = EAHere :F-P_ą at any x dF ! → At every section unth only arial load do Lo EA EA PLo → Elongation = A 3.144/4 → Elongation = 3.Here :F-P_ą at any x dF do → At every section uith only arial load 1 Lo EA EA PLo → Elongation = A- 3.14 d^/4 4 P Lo